### v2

@@ 1 Introduction
Using this theory one can define, on a metric graph $\Gamma$, the gonality $\operatorname{gon}(\Gamma)$ (see Definition2) and the Clifford Index $\operatorname{Cliff}(\Gamma)$ (see Definition 4) in the same way as in the case of a smooth curve $C$. As in the case of smooth curves the inequality $\operatorname{Cliff}(\Gamma)\leq\operatorname{gon}(\Gamma)-2$ holds by definition. In the case of a smooth curve $C$ the inequality $\operatorname{gon}(C)-3\leq\operatorname{Cliff}(C)$ is proved in [ref2]. Whether or not this inequality also holds for metric graphs is mentioned as an unknown fact in [ref3]. The importance of this inequality comes from the fact that $\operatorname{gon}(\Gamma)$ is defined using divisors of rank 1 on $\Gamma$, while $\operatorname{Cliff}(\Gamma)$ gives information on divisors on $\Gamma$ of any prescribed rank.
@@
Chains of cycles were considered for the first time in [ref4] where the authors used those metric graphs to obtain a tropical proof of the Brill-Noether Theorem. Later on those metric graphs are used to prove many results in Brill-Noether Theory (see some applications in [ref1], see e.g. [ref5]).
@@
In the second section we give a short overview on the results of [ref7] and we make the statement on $\underline{m}$-displacement tableaux. In the third section we give a proof of that statement.
@@ 2 Some results on divisor theory on chains of cycles
The rank $r=\operatorname{rk}(D)$ of a divisor $D$ on $\Gamma$ is the largest integer $n$ such that for each effective divisor $E$ of degree $n$ on $\Gamma$ there exists an effective divisor $D^{\prime}$ on $\Gamma$ equivalent to $D$ containing $E$. (In case $D$ is not equivalent to an effective divisor then we put $\operatorname{rk}(D)=-1$.)
@@
The gonality $n=\operatorname{gon}(\Gamma)$ of $\Gamma$ is the smallest integer $m$ such that there exists a divisor $D$ on $\Gamma$ of degree $m$ such that $\operatorname{rk}(D)\geq 1$.
On a metric graph $\Gamma$ of genus $g$ we have a canonical divisor $K_{\Gamma}$ and there is a Riemann-Roch Theorem as in the case of curves. Divisors having rank at most 0 are not interesting, so we are only interested in divisors $D$ satisfying $\operatorname{rk}(D)\geq 1$ and $\operatorname{rk}(K_{\Gamma}-D)\geq 1$. Because of the Riemann-Roch Theorem this second condition is equivalent to $\operatorname{rk}(D)-d+g\geq 2$.
@@
The Clifford index $c=\operatorname{Cliff}(\Gamma)$ is the smallest integer $m$ such that there exists a divisor $D$ with $\operatorname{rk}(D)\geq 1$ and $\operatorname{rk}(D)-d+g\geq 2$ with $\operatorname{Cliff}(D)=m$.
@@
The length of those segments does not influence the divisor theory on $\Gamma$. In particular it can be taken equal to zero. Taking it nonzero is handsome to distinguish e.g. the points $w_{1}$ on $C_{1}$ and $v_{2}$ on $C_{2}$. In order to prove Theorem 1 it is enough to prove the following statement on such a graph $\Gamma$. If $D$ is a divisor of degree $d$ and rank at least $r$ on $\Gamma$ with $r\geq 1$ and $r-d+g\geq 2$ then there exists a divisor $D^{\prime}$ of degree $d-2r+2$ and rank at least 1 on $\Gamma$.
@@
For positive integers $m$ and $n$ we write $[m\times n]$ to denote the set $\{1,\cdots,m\}\times\{1,\cdots,n\}$. An $\underline{m}$-displacement tableau on $[m\times n]$ of genus $g$ is a function $t:[m\times n]\rightarrow\{1,\cdots,g\}$ such that
@@
Given $r\in\mathbb{Z}_{\geq 1}$ and $d\in\mathbb{Z}$ with $r-d+g\geq 2$. There exists a divisor $D$ on $\Gamma$ of degree $d$ and rank at least $r$ if and only if there exists an $\underline{m}$-displacement tableau $t:[(g-d+r)\times(r+1)]\rightarrow\{1,\cdots,g\}$.
So in order to prove Theorem 1 we need to prove
@@
Let $r\in\mathbb{Z}_{\geq 1}$ and $d\in\mathbb{Z}$ with $r-d+g\geq 2$. If there exists an $\underline{m}$-displacement tableau $t:[(g-d+r)\times(r+1)]\rightarrow\{1,\cdots,g\}$ then there exists an $\underline{m}$-displacement tableau $t^{\prime}:[(g-d+2r-1)\times 2]\rightarrow\{1,\cdots,g\}$.
@@ 3 Proof of the theorem
Lemma 1.
The statement of Theorem 2 holds for $d\geq g$.
Proof.
Assume $d\geq g$. Then $K_{\Gamma}-D$ is a divisor of degree $d^{\prime}=2g-2-d$ and rank at least equal to $r^{\prime}=r+g-d-1$. Since $g-d+r\geq 2$ we have $r^{\prime}\geq 1$ while $g-d^{\prime}+r^{\prime}=r+1\geq 2$. Moreover $r^{\prime}\leq r-1$, so we can apply the induction hypothesis and conclude that there is an $\underline{m}$-displacement tableau on $[(g-d^{\prime}+2r^{\prime}-1)\times 2]=[(g-d+2r-1)\times 2]$. ∎
So from now on in the proof we assume $d\leq g-1$.
In the proof we are going to use tableaux associated to ”smaller” chains of cycles. We use the following terminology. If $11\leq k\leq g-$ 1 then by $C_{1}\cup\cdots\cup C_{k}$ we mean the chain of cycles obtained from $\Gamma$ by deleting all cycles $C_{k+1}\cup\cdots\cup C_{g}$ and the segments from $w_{i}$ to $v_{i+1}$ for $k\leq i\leq g-1$. Its torsion profile is given by $(m_{2},\cdots,m_{k})$; we still talk about $\underline{m}$-displacement tableaux on $C_{1}\cup\cdots\cup C_{k}$ (so we use the same notation to indicate the torsion profile).
Let $g^{\prime}=t(g-d+r,r)$, hence $g^{\prime}\leq g-1$. Then $t$ restricted to $[(g-d+r)\times r]=(g^{\prime}-(d-(g-g^{\prime})-1)+(r-1))\times((r-1)+1)]$ is an $\underline{m}$-displacement tableau for $C_{1}\cup\cdots\cup C_{g^{\prime}}$. Since $r\geq 2$ of course $r-1\geq 1$ while $g^{\prime}-(d-(g-g^{\prime})-1)+(r-1)=g-d+r\geq 2$. From the induction hypothesis we obtain an $\underline{m}$-displacement tableau $\tilde{t^{\prime}}$ on $[(g^{\prime}-(d-(g-g^{\prime})-1)+2(r-1)-1)\times 2]=[(g-d+2r-2)\times 2]$ for $C_{1}\cup\cdots\cup C_{g^{\prime}}$. In case $g^{\prime}\leq g-2$ then we enlarge $\tilde{t^{\prime}}$ with $t^{\prime}(g-d+2r-1,1)=g-1$ and $t^{\prime}(g-d+2r-1,2)=g$ and we obtain an $\underline{m}$-displacement tableau on $[(g-d+2r-1)\times 2]$ for $\Gamma$.
To continue the proof we can assume that $g^{\prime}=t(g-d+r,r)=g-1$, hence also $(g-d+r,r+1)=g$. In case $t(g-d+r-1,r+1)=g-1$ then $m_{g-1}=2$ and by extending $\tilde{t^{\prime}}$ with $t^{\prime}(g-d+2r-1,1)=g-1$ and $t^{\prime}(g-d+2r-1,2)=g$ we obtain an $\underline{m}$-displacement tableau on $[(g-d+2r-1)\times 2]$ for $\Gamma$.
To continue the proof we can assume that $t(g-d+r-1,r+1)\leq g-2$. It implies $g^{\prime}=t(g-d+r-1,r)\leq g-3$ (so we changed the meaning of the symbol $g^{\prime}$), so the restriction of $t$ to $[(g-d+r-1)\times r]=[(g^{\prime}-(d-(g-g^{\prime}))+(r-1)\times((r-1)+1)]$ is an $\underline{m}$-displacement tableau for $C_{1}\cup\cdots\cup C_{g^{\prime}}$ with $r-1\geq 1$ and $g^{\prime}-(d-(g-g^{\prime}))+(r-1)=g-d+r-1$. In case $g-d+r=2$ then from $d\leq g-1$ we obtain $r\leq 1$ contradicting our assumption $r\geq 2$. So we find $g-d+r-1\geq 2$ and we can use the induction hypothesis giving the existence of an $\underline{m}$-displacement tableau $\tilde{t^{\prime}}$ on $[(g^{\prime}-(d-(g-g^{\prime}))+2(r-1)-1)\times 2]=[(g-d+2r-3)\times 2]$ on $C_{1}\cup\cdots\cup C_{g^{\prime}}$. In case $g^{\prime}\leq g-4$ we can extend $\tilde{t^{\prime}}$ putting $t^{\prime}(g-d+2r-2,1)=g-3$, $t^{\prime}(g-d+2r-2,2)=g-2$, $t^{\prime}(g-d+2r-1,1)=g-1$ and $t^{\prime}(g-d+2r-1,2)=g$. We obtain an $\underline{m}$-displacement tableau on $[(g-d+2r-1)\times 2]$ for $\Gamma$.
To continue the proof we can assume $t(g-d+r-1,r)=g-3$ and hence also $t(g-d+r-1,r+1)=g-2$.
Now we consider the following claim involving a positive integer $k$.
Claim.
Let $k$ be an integer satisfying $11\leq k\leq g-d+r-$ 1. If there exist some $0\leq k^{\prime}\leq k$ such that $t(g-d+r-k^{\prime},r+1)\neq g-2k^{\prime}$ or $t(g-d+r-k^{\prime},r)\neq g-2k^{\prime}-1$ then $\Gamma$ has an $\underline{m}$-displacement tableau on $[(g-d+2r-1)\times 2]$.
In case Theorem 2 would not be true then the claim implies $g-2k\geq g-d+2r-k$, i.e. $k\leq d-2r$. Also in case $k=d-2r$ it would imply $t(i,j)=i+j-1$ for $1\leq i\leq g-d+r-k$ and $11\leq j\leq r+$ 1. In the above arguments we proved the claim for $k=1$. In case $k<d-2r$ and $k<g-d+r-1$ we are going to prove that the claim implies the similar claim replacing $k$ by $k+1$. In the other cases we are going to prove that the claim implies the existence of an $\underline{m}$-replacement tableau on $[(g-d+2r-1)\times 2]$. In this way the proof of Theorem 2 becomes complete.
First we assume the claim holds for some value $k$ satisfying $k<d-2r$ and $g-d+r-k>1$ and we assume no value $k^{\prime}$ as in the claim exists. Restricting $t$ to $[(g-d+r-k)\times r]=[((g-2k-1)-(d-k-2)+(r-1))\times((r-1)+1)]$ we obtain an $\underline{m}$-displacement tableau on $C_{1}\cup\cdots\cup C_{g-2k-1}$. We have $r-1\geq 1$ and $(g-2k-1)-(d-k-2)+(r-1)=g-k-d+r\geq 2$. By assumption we obtain an $\underline{m}$-tableau $\tilde{t^{\prime}}$ on $[((g-2k-1)-(d-k-2)+2(r-1)-1)\times 2]=[(g-d+2r-k-2)\times 2]$ for $C_{1}\cup\cdots\cup C_{g-2k-1}$. In case $t(g-d+r-k-1,r+1)=g-2k-1$ we have $m_{g-2k-1}=2$. Then we can extend $\tilde{t^{\prime}}$ by $t^{\prime}(g-d+2r-k-2+x,1)=g-2k-3+2x$ and $t^{\prime}(g-d+2r-k-2+x,2)=g-2k-2+2x$ for $11\leq x\leq k+$ 1, which gives rise to an $\underline{m}$-displacement tableau of $\Gamma$.
To continue the proof we can assume $t(g-d+r-k-1,r+1)\leq g-2k-2$. In case $t(g-d+r-k-1,r)\leq g-2k-4$ then the restriction of $t$ to $[(g-d+r-k-1)\times r]=[((g-2k-4)-(d-k-4)+(r-1))\times((r-1)+1)]$ is an $\underline{m}$-displacement tableau for $C_{1}\cup\cdots\cup C_{g-2k-4}$. We have $r-1\geq 1$ and $(g-2k-4)-(d-k-4)+(r-1)=g-d-k+r-1\geq 1$.
In case $g-d-k+r=2$ then the restriction of $t$ to $[(g-d+r-k)\times(r+1)]=[((g-2k)-(d-k)+r)\times(r+1)]$ is an $\underline{m}$-displacement tableau on $C_{1}\cup\cdots\cup C_{g-2k}$. We have $(g-2k)-(d-k)+r=g-d-k+r\geq 2$. In case $d-k\leq(g-2k)-1$ we have $g\geq d+k+1$, so $d+k+2-r\geq d+k+1$ and therefore $r\leq 1$, contradicting our assumption. So we find $d-k\geq g-2k$ and then Lemma 1 implies the existence of an $\underline{m}$-displacement tableau $\tilde{t^{\prime}}$ on $[((g-2k)-(d-k)+2r-1)\times 2]=[(g-k-d+2r-1)\times 2]$ for $C_{1}\cup\cdots\cup C_{g-2k}$. Using the remaining $2k$ values $\{g-2k+1,\cdots,g\}$ we can complete $\tilde{t^{\prime}}$ to an $\underline{m}$-displacement tableau on $[(g-d+2r-1)\times 2]$ for $\Gamma$.
@@
In case $g-d-k+r\geq 3$ then by assumption we find an $\underline{m}$-displacement tableau $\tilde{t^{\prime}}$ on $[((g-2k-4)-(d-k-4)+2(r-1)-1)\times 2]=[(g-k-d+2r-3)\times 2]$ on $C_{1}\cup\cdots\cup C_{g-2k-4}$. We can extend $\tilde{t^{\prime}}$ by $t^{\prime}(g-k-d+2r-3+i,1)=g-2k-4+2i-1$ and $t^{\prime}(g-k-d+2r-3+i,2)=g-2k-4+2i$ for $21\leq i\leq k+$ 2 and we obtain an $\underline{m}$-displacement tableau on $[(g-d+2r-1)\times 2]$ on $\Gamma$.
To continu the proof we can assume $t(g-d+r-k-1,r)=g-2k-3$ and therefore also $t(g-d+r-k-1,r+1)=g-2k-2$. So we obtain the claim with $k$ replaced by $k+1$.
Next we assume the claim holds for some value $k$ with $g-d+r-k>1$ but $k=d-2r$. As remarked we have $t(i,j)=i+j-1$ for $1\leq i\leq g-d+r-k$ and $11\leq j\leq r+$ 1. It follows $m_{i}=2$ for $12\leq i\leq g-d+2r-k-$ 1. So $\tilde{t^{\prime}}(i,j)=i+j-1$ is an $\underline{m}$-displacement tableau on $[(g-d+2r-k-1)\times 2]$ for $C_{1}\cup\cdots\cup C_{g-d+2r-k}$. But the $d+k-2r=2k$ values from $\{g-d+2r-k+1,\cdots,g\}$ can be used to extend $\tilde{t^{\prime}}$ to an $\underline{m}$-displacement tableau on $[(g-d+2r-1)\times 2]$ for $\Gamma$.
Finally we assume $g-d+r-k=1$. Then the restriction of $t$ to $[2\times(r+1)]=[((g-2k+2)-(g+r-2k)+r)\times(r+1)]$ is an $\underline{m}$-displacement tableau on $C_{1}\cup\cdots\cup C_{g-2k+2}$. Note that $g+r-2k\geq 0$ since $g-2k\geq 0$. Also $(g-2k+2)-(g+r-2k)+r=2$ but $r+g-2k\geq g-2k+2$. So from Lemma 1 we obtain there exists an $\underline{m}$-displacement tableau $\tilde{t^{\prime}}$ on $[((g-2k+2)-(g+r-2k)+2r-1)\times 2]=[(r+1)\times 2]$ for $C_{1}\cup\cdots\cup C_{g-2k+2}$. The set $\{g-2k+3,\cdots,g\}$ consists of $22k-$ 2 unused values in $\tilde{t^{\prime}}$. Since $(r+1)+(k-1)=(d-r+k+1)-d+2r-1=g-d+2r-1$ those values can be used to obtain an $\underline{m}$-displacement tableau on $[(g-d+2r-1)\times 2]$ for $\Gamma$.
This finishes the proof of Theorem 2, hence also of Theorem 1.

### v3

@@ 1 Introduction
Using this theory, one can define, on a metric graph $\Gamma$, the gonality $\operatorname{gon}(\Gamma)$ (see Definition 2) and the Clifford index $\operatorname{Cliff}(\Gamma)$ (see Definition 4) in the same way as in the case of a smooth curve $C$. As in the case of smooth curves the inequality $\operatorname{Cliff}(\Gamma)\leq\operatorname{gon}(\Gamma)-2$ holds by definition. In the case of a smooth curve $C$ the inequality $\operatorname{gon}(C)-3\leq\operatorname{Cliff}(C)$ is proved in [ref2]. Whether or not this inequality also holds for metric graphs is mentioned as an unknown fact in [ref3]. The importance of this inequality comes from the fact that $\operatorname{gon}(\Gamma)$ is defined using divisors of rank 1 on $\Gamma$, while $\operatorname{Cliff}(\Gamma)$ gives information on divisors on $\Gamma$ of any prescribed rank.
@@
Chains of cycles were considered for the first time in [ref4] where the authors used those metric graphs to obtain a tropical proof of the Brill–Noether Theorem. Later on those metric graphs are used to prove many results in Brill–Noether Theory (see some applications in [ref1], see e.g. [ref5]).
@@
In the second section we give a short overview on the results of [ref7] and we make a statement (see Theorem 2) on $\underline{m}$-displacement tableaux implying Theorem 1. In the third section we give a proof of that statement.
@@ 2 Some results on divisor theory on chains of cycles
The rank $\operatorname{rk}(D)$ of a divisor $D$ on $\Gamma$ is the largest integer $r$ such that for each effective divisor $E$ of degree $r$ on $\Gamma$ there exists an effective divisor $D^{\prime}$ on $\Gamma$ equivalent to $D$ containing $E$. (In case $D$ is not equivalent to an effective divisor then we put $\operatorname{rk}(D)=-1$.)
@@
The gonality $\operatorname{gon}(\Gamma)$ of $\Gamma$ is the smallest integer $d$ such that there exists a divisor $D$ on $\Gamma$ of degree $d$ such that $\operatorname{rk}(D)\geq 1$.
On a metric graph $\Gamma$ of genus $g$ we have a canonical divisor $K_{\Gamma}$ and there is a Riemann–Roch Theorem as in the case of curves. Divisors having rank at most 0 are not interesting, so we are only interested in divisors $D$ satisfying $\operatorname{rk}(D)\geq 1$ and $\operatorname{rk}(K_{\Gamma}-D)\geq 1$. Because of the Riemann-Roch Theorem this second condition is equivalent to $\operatorname{rk}(D)-d+g\geq 2$.
@@
The Clifford index $\operatorname{Cliff}(\Gamma)$ is the smallest integer $c$ such that there exists a divisor $D$ with $\operatorname{rk}(D)\geq 1$ and $\operatorname{rk}(D)-d+g\geq 2$ with $\operatorname{Cliff}(D)=c$.
@@
The length of those segments does not influence the divisor theory on $\Gamma$. In particular it can be taken equal to zero. Taking it non-zero is handsome to distinguish e.g. the points $w_{1}$ on $C_{1}$ and $v_{2}$ on $C_{2}$. In order to prove Theorem 1 it is enough to prove the following statement on such a graph $\Gamma$. If $D$ is a divisor of degree $d$ and rank at least $r$ on $\Gamma$ with $r\geq 1$ and $r-d+g\geq 2$ then there exists a divisor $D^{\prime}$ of degree $d-2r+2$ and rank at least 1 on $\Gamma$.
@@
For positive integers $m$ and $n$ we write $[m\times n]$ to denote the set $\{1,\cdots,m\}\times\{1,\cdots,n\}$. Let $\underline{m}=(m_{2},\cdots,m_{g})$ be a sequence of $g-1$ non-negative integers. An $\underline{m}$-displacement tableau on $[m\times n]$ is a function $t:[m\times n]\rightarrow\{1,\cdots,g\}$ such that
@@
Let $\Gamma$ be a chain of cycles of genus $g$ with torsion profile $\underline{m}$. Given $r\in\mathbb{Z}_{\geq 1}$ and $d\in\mathbb{Z}$ with $r-d+g\geq 2$. There exists a divisor $D$ on $\Gamma$ of degree $d$ and rank at least $r$ if and only if there exists an $\underline{m}$-displacement tableau $t:[(g-d+r)\times(r+1)]\rightarrow\{1,\cdots,g\}$.
So in order to prove Theorem 1 it is enough to prove
@@
Let $\underline{m}=(m_{2},\cdots,m_{g})$ be a sequence of $g-1$ non-negative integers. Let $r\in\mathbb{Z}_{\geq 1}$ and $d\in\mathbb{Z}$ with $r-d+g\geq 2$. If there exists an $\underline{m}$-displacement tableau $t:[(g-d+r)\times(r+1)]\rightarrow\{1,\cdots,g\}$ then there exists an $\underline{m}$-displacement tableau $t^{\prime}:[(g-d+2r-1)\times 2]\rightarrow\{1,\cdots,g\}$.
@@ 3 Proof of the theorem
In the proof we are going to use tableaux obtained from restricting $t$ to subrectangles of $[(g-d+r)\times(r+1)]$. They are displacement tableaux using subsequences of $\underline{m}$. Therefore, for $1\leq j\leq g$ we introduce $\underline{m}_{j}=(m_{2},\cdots,m_{j})$ (for $j=g$ this is $\underline{m}$, for $j=1$ this is empty). In particular an $\underline{m}_{j}$-displacement tableau has values in $\{1,\cdots,j\}$.
@@
Since $t$ is an $\underline{m}$-displacement tableau we have $g-d+2r\leq g$, equivalently $d\geq 2r$. In case $d=2r$ we have $t(i,j)=i+j-1$ on $[(g-d+r)\times(r+1)]$. It follows that $m_{2}=\cdots=m_{g-1}=2$. Then on $[(g-d+2r-1)\times 2]=[(g-1)\times 2]$ we find $t^{\prime}(i,j)=i+j-1$ is an $\underline{m}$-displacement tableau.
So, from now on, we can assume $d>2r$.
Step $\underline{\text{Step 1}}$: Theorem 2 holds if $t(g-d+r,r)<g-1$.
If $t(g-d+r,r)<g-1$ then the restriction of $t$ to $[(g-d+r)\times r]=[((g-2)-(d-3)+(r-1))\times((r-1)+1)]$ is an $\underline{m}_{g-2}$-displacement tableau. Since $r\geq 2$ of course $r-1\geq 1$ while $(g-2)-(d-3)+(r-1)=g-d+r\geq 2$, so we can apply the induction hypothesis to that restriction. We obtain the existence of an $\underline{m}_{g-2}$-tableau $\tilde{t^{\prime}}$ on $[((g-2)-(d-3)+2(r-1)-1)\times 2]=[(g-d+2r-2)\times 2]$. We obtain an $\underline{m}$-displacement tableau $t^{\prime}$ on $[(g-d+2r-1)\times 2]$ having restriction $\tilde{t^{\prime}}$ to $[(g-d+2r-2)\times 2]$ and taking $t^{\prime}(g-d+2r-1,1)=g-1$ and $t^{\prime}(g-d+2r-1,2)=g$ (further on we will say $\tilde{t^{\prime}}$ is extended).
In case $t(g-d+r,r)\geq g-1$ then $t(g-d+r,r)=g-1$ and $t(g-d+r,r+1)=g$. Because of Step 1, from now on we can assume those two equalities hold.
Step $\underline{\text{Step 2}}$: Theorem 2 holds if $t(g-d+r-1,r+1)=g-1$.
If $t(g-d+r-1,r+1)=g-1=t(g-d+r,r)$ then $m_{g-1}=2$. The restriction of $t$ to $[(g-d+r)\times r]=[((g-1)-(d-2)+(r-1))\times((r-1)+1)]$ is an $\underline{m}_{g-1}$-displacement tableau. Again we can apply the induction hypothesis implying there exists an $\underline{m}_{g-1}$-tableau $\tilde{t^{\prime}}$ on $[((g-1)-(d-2)+2(r-1)-1)\times 2]=[(g-d+2r-2)\times 2]$. Because $m_{g-1}=2$ we can extend it to an $\underline{m}$-displacement tableau $t^{\prime}$ on $[(g-d+2r-1)\times 2]$ taking $t^{\prime}(g-d+2r-1,1)=g-1$ and $t^{\prime}(g-d+2r-1,2)=g$.
From now on we can assume also $t(g-d+r-1,r+1)\leq g-2$ (of course $t(g-d+r-1,r+1)>g-1$ is not possible).
This inequality implies $g-d+2r-1\leq g-2$, equivalently $d\geq 2r+1$. In case $d=2r+1$ we obtain $t(i,j)=i+j-1$ for $(i,j)\in[(g-d+r-1)\times(r+1)]$. It follows $m_{2}=\cdots=m_{g-3}=2$. We find an $\underline{m}$-displacement tableau $t^{\prime}$ on $[(g-d+2r-1)\times 2]=[(g-2)\times 2]$ by taking $t^{\prime}(i,j)=i+j-1$ in case $(i,j)\in[(g-3)\times 2]$, $t^{\prime}(g-2,1)=g-1$ and $t^{\prime}(g-2,2)=g$. Therefore from now on we can also assume $d>2r+1$.
Step $\underline{\text{Step 3}}$: Theorem 2 holds if $t(g-d+r-1,r)<g-3$.
If $t(g-d+r-1,r)<g-3$ then the restriction of $t$ to $[(g-d+r-1)\times r]=[((g-4)-(d-4)+(r-1))\times((r-1)+1)]$ is an $\underline{m}_{g-4}$-displacement tableau. Since $r\geq 2$ we still have $r-1\geq 1$. We have $(g-4)-(d-4)+(r-1)=g-d+r-1$. In case $g-d+r=2$ we have that $t^{\prime}(i,j)=t(j,i)$ defines an $\underline{m}$-displacement tableau on $[(r+1)\times 2]=[(2r-r+1)\times 2]=[(2r+g-d-1)\times 2]$, so the theorem is proved in that case. In case $g-d+r>2$ we have $(g-4)-(d-4)+(r-1)\geq 2$ and we can use the induction hypothesis. In that case we obtain the existence of an $\underline{m}_{g-4}$-displacement tableau $\tilde{t^{\prime}}$ on $[((g-4)-(d-4)+2(r-1)-1)\times 2]=[(g-d+2r-3)\times 2]$. This can be extended to an $\underline{m}$-displacement tableau $t^{\prime}$ on $[(g-d+2r-1)\times 2]$ by taking $t^{\prime}(g-d+2r-2,1)=g-3$, $t^{\prime}(g-d+2r-2,2)=g-2$, $t^{\prime}(g-d+2r-1,1)=g-1$ and $t^{\prime}(g-d+2r-1,2)=g$.
So we proved Theorem 2 unless $t(g-d+r-1,r+1)\leq g-2$ and $t(g-d+r-1,r)\geq g-3$. From the definition of a tableau those inequalities are equivalent to $t(g-d+r-1,r+1)=g-2$ and $t(g-d+r-1,r)=g-3$. From now on we assume those equalities.
For an integer $11\leq k\leq g-d+r-$ 1 we formulate the following assumption.
$\underline{\text{Assumption(k)}}$: For all integers $0\leq k^{\prime}\leq k$ we have $t(g-d+r-k^{\prime},r+1)=g-2k^{\prime}$ and $t(g-d+r-k^{\prime},r)=g-2k^{\prime}-1$.
From the previous steps it follows we can assume that Assumption(1) holds. Also in case Assumption(k) holds then we need $g-2k\leq g-d+2r-k$, or equivalently $d\leq 2r+k$. In case $d=2r+k$ we have $t(i,j)=i+j-1$ for all $(i,j)\in[(g-d+r-k)\times(r+1)]$ In particular $m_{2}=\cdots=m_{g-2k-1}=2$.
Now we make an induction step on $k$.
Step $\underline{\text{Step 4}}$: Assume $k<g-d+r-1$ and $d>2r+k$. If Assumption(k) holds but Assumption(k+1) does not hold then Theorem 2 holds.
The proof of this induction step is similar to steps 2 and 3.
Restricting $t$ to $[(g-d+r-k)\times r]=[((g-2k-1)-(d-k-2)+(r-1))\times((r-1)+1)]$ we obtain an $\underline{m}_{g-2k-1}$-displacement tableau. We have $r-1\geq 1$ and $(g-2k-1)-(d-k-2)+(r-1)=g-k-d+r\geq 2$. By assumption we obtain an $\underline{m}_{g-2k-1}$-tableau $\tilde{t^{\prime}}$ on $[((g-2k-1)-(d-k-2)+2(r-1)-1)\times 2]=[(g-d+2r-k-2)\times 2]$. In case $t(g-d+r-k-1,r+1)=g-2k-1$ we have $m_{g-2k-1}=2$. Then we can extend $\tilde{t^{\prime}}$ by $t^{\prime}(g-d+2r-k-2+x,1)=g-2k-3+2x$ and $t^{\prime}(g-d+2r-k-2+x,2)=g-2k-2+2x$ for $11\leq x\leq k+$ 1, which gives rise to an $\underline{m}$-displacement tableau on $[(g-d+2r-1)\times 2]$.
To continue the proof we can assume $t(g-d+r-k-1,r+1)\leq g-2k-2$. In case $t(g-d+r-k-1,r)\leq g-2k-4$ then the restriction of $t$ to $[(g-d+r-k-1)\times r]=[((g-2k-4)-(d-k-4)+(r-1))\times((r-1)+1)]$ is an $\underline{m}_{g-2k-4}$-displacement tableau. We have $r-1\geq 1$ and $(g-2k-4)-(d-k-4)+(r-1)=g-d-k+r-1\geq 1$.
In case $g-d-k+r=2$ then we have $t(g-d+r-k,r+1)=t(2,r+1)=g-2k$. We consider the $\underline{m}_{g-2k}$-displacement tableau $\tilde{t^{\prime}}$ on $[(r+1)\times 2]=[(2r+1-(d-g+k+2))\times 2]=[(g-d+2r-1-k)\times 2]$ given by $\tilde{t^{\prime}}(i,j)=t(j,i)$. This can be extended to an $\underline{m}$-displacement tableau $t^{\prime}$ on $[(g-d+2r-1)\times 2]$ by taking $t^{\prime}(g-d+2r-1-i,1)=g-2i-1$ and $t^{\prime}(g-d+2r-1-i,2)=g-2i$ for $10\leq i\leq k-$ 1.
In case $g-d-k+r>2$ then the induction hypothesis gives the existence of an $\underline{m}_{g-2k-4}$-displacement tableau $\tilde{t^{\prime}}$ on $[((g-2k-4)-(d-k-4)+2(r-1)-1)\times 2]=[(g-d-k+2r-3)\times 2]$. This can be extended to an $\underline{m}$-displacement tableau $t^{\prime}$ on $[(g-d+2r-1)\times 2]$ putting $t^{\prime}(g-d+2r-k-2+x,1)=g-2k-3+2x$ and $t^{\prime}(g-d+2r-k-2+x,2)=g-2k-2+2x$ for $10\leq x\leq k+$ 1.
So we proved Theorem 2 unless $t(g-d+r-k-1,r)\geq g-2k-3$ and $t(g-d+r-k-1,r+1)\leq g-2k-2$. By definition of a tableau we find equalities in this exceptional case and therefore Assumption(k+1).
So we have proved Theorem 2 unless there exists some integer $11\leq k\leq g-d+r-$ 1 such that Assumption(k) holds with $d=2r+k$ or $k=g-d+r-1$.
Step $\underline{\text{Step 5}}$: If Assumption(k) holds with $k=d-2r$ then Theorem 2 holds.
It is already noted that in this case $m_{2}=\cdots=m_{g-d+2r-k-1}=2$. So $\tilde{t^{\prime}}(i,j)=i+j-1$ defines an $\underline{m}_{g-d+2r-k}$-displacement tableau on $[(g-d+2r-1-k)\times 2]$. This can be extended to become an $\underline{m}$-displacement tableau $t^{\prime}$ on $[(g-d+2r-1)\times 2]$ taking $t^{\prime}(g-d+2r-1-i,1)=g-2i-1$ and $t^{\prime}(g-d+2r-1-i,2)=g-2i$ for $10\leq i\leq k-$ 1.
Step $\underline{\text{Step 6}}$: If Assumption(k) holds with $g-d+r-k=1$ then Theorem 2 holds.
Since $t(g-d+r-(g-d+r-2),r+1)=t(2,r+1)=g-2(k-1)$ we obtain an $\underline{m}_{g-2(k-1)}$-displacement tableau $\tilde{t^{\prime}}$ on $[(r+1)\times 2]$. But $r+1=2r+1-(d+k-g+1)=g-d+2r-k$ so we can extend $\tilde{t^{\prime}}$ to an $\underline{m}$-displacement tableau $t^{\prime}$ on $[(g-d+2r-1)\times 2]$ taking $t^{\prime}(g-d+2r-1-i,1)=g-2i-1$ and $t^{\prime}(g-d+2r-1-i,2)=g-2i$ for $20\leq i\leq k-$ 2.
This finishes the proof of Theorem 2.
Acknowledgement. I am grateful to the referee for his suggestions to improve the paper.